Prove that Homz(Z,M) is isomorphic to M where M is an abelian group.


1
Expert's answer
2021-06-15T12:32:14-0400

Define a function "\phi:\text{Hom}(\mathbb{Z},G) \to G"

such that "\phi(f)=f(1) \forall f \in \text{Hom}(\mathbb{Z},G)."

This function is well-defined since "f(n) = [f(1)]^n \forall n \in \mathbb{Z}."

Let "f,g \in \text{Hom}(\mathbb{Z},G)." Consider

"\phi(fg)=(fg)(1)=f(1)g(1)=\phi(f)\phi(g)."

Hence "\phi" is a homomorphism.

Next, suppose "\phi(f)=\phi(g)." Then "f(1)=g(1)." So "\forall n \in \mathbb{Z}, f(n) = [f(1)]^n\\ = [g(1)]^n=g(n)."

Thus "f = g." Hence "\phi" is one-to-one.

Finally, let "x \in G." Define a function "\psi:\mathbb{Z}\to G" such that "\psi(n)=x^n\forall n \in \mathbb{Z}." Let "m,n \in \mathbb{Z}." Consider

"\psi(m+n) = x^{m+n}=x^mx^n = \psi(m)\psi(n)."

So "\psi" is a homomorphism.

Hence "\phi(\psi) = \psi(1)=x." Thus "\psi" is onto.

Hence "\phi" is an isomorphism.